OpenStudy (anonymous):

The formula, A = P + Prt, can be used to find the amount of money (A) in a savings account, where P is the initial deposit, r is the interest rate, and t is the time in years. Jonathan deposits $400 into a new savings account with an interest rate of 3%. He makes no other deposits or withdrawals. How much money, in dollars, will Jonathan have in the account after 1 year?

OpenStudy (anonymous):

A = 400 + 400(0.03)(1) A = 400 + 12 A = $412